# Non-linear degenerate parabolic flow equations and a finer differential structure on Wasserstein spaces

### v1

@@ Abstract
We define new differential structures on the Wasserstein spaces $\mathcal{W}_{p}(M)$ for $p>2$ and a general Riemannian manifold $(M,g)$. We consider a very general and possibly degenerate second order partial differential flow equation with measure dependent coefficients to expand the notion of smooth curves and to ensure that the new differential structure is finer than the classical one. Under weak assumptions, we explicitly construct smooth solutions as uniform limits of Average Flow Approximation Series (a variant of explicit Euler–scheme approximations) in $\mathcal{W}_{p}(M)$ and, thus, prove a generalzed version of the Central Limit Theorem. Under slightly stronger assumptions, we prove that smooth solutions of our newly introduced flow–equation are unique.
@@ Acknowledgements
I am grateful to Alessio Figalli and Ulisse Stefanelli for their guidance and support. I would also like to thank Hans Gersbach, Wilhelm Schlag, Darko Mitrovic, Nikolai V. Krylov and Hermann Schichl for their valuable input. All errors are mine.
@@ 1 Introduction
Recent years have seen an increased need for analytical calculus on spaces other than finite dimensional vector spaces. In particular, measure spaces like the Wasserstein spaces of optimal transport theory have been of central interest in prominent works like Ambrosio et al. [1, 2], Figalli & Glaudo [7], Santambrogio [21], Villani et al. [23], Villani [22].
@@
on the Wasserstein space $\mathcal{W}_{p}(M)$, for a Riemannian manifold $(M,g)$, $p\geq 1$ and some probability vector–field $V$, which remains to be defined. To these means, Otto & Villani [18], Otto [17] were the first to endow the Wasserstein spaces $\mathcal{W}_{p}(\mathbb{R}^{n})$ for $p>1$ with differential structures. The so-called continuity equation
@@
Detailed expositions of the theory in $\mathbb{R}^{n}$ and more general spaces can be found in Ambrosio et al. [1], Gigli [9], Bonnet & Frankowska [4], Cavagnari et al. [5]. However, the so–defined differential structures have important limitations.
@@ 2 Preliminaries
In this preliminary chapter, we introduce some basic notations we shall use in this work and recall some classical results in Optimal transport theory and Riemannian geometry. Furthermore, we prove some simple results, which shall be of use later.
2.1 Notations
@@
We will need some real function spaces, especially Sobolev spaces, throughout this work. Let $(M,g)$ be a Riemannian manifold and $A\subset M$. We denote by $\lambda$ the Lebesgue measure on $\mathbb{R}^{n}$ and define the function spaces (see Michor [14] for formal introductions):
$\mathcal{W}_{p}(M)$ The Wasserstein space of pp–integrable probability measures on $M$ for $p\geq 1$ (see [7]).
$\mathcal{W}_{p}(\mu,\nu)$ The pp–Wasserstein distance between $\mu,\nu\in\mathcal{W}_{p}(M)$, for $p\geq 1$ (see [7]).
@@
$C(A,\mathbb{R}^{m})$ The set of all continuous functions $f:\mathbb{R}^{n}\supseteq A\to\mathbb{R}^{m}$.
$C^{k}(A,\mathbb{R}^{m})$ For $k\in\mathbb{N}\cup\{\infty,\omega\}$ the set of all kk–times continuously differentiable functions $f\in C(A,\mathbb{R})$. For $k=\infty$ the set of all smooth functions, and for $k=\omega$ the set of all real analytic functions. Also, $C^{0}(A,\mathbb{R}^{m}):=C(A,\mathbb{R}^{m})$.
$C^{k}_{c}(A,\mathbb{R}^{m})$ For $k\in\mathbb{N}\cup\{\infty,\omega\}$ the set of all $f\in C^{k}(A,\mathbb{R})$ with ${\operatorname{supp}}f$ compact.
$C^{k}_{b}(A,\mathbb{R}^{m})$ For $k\in\mathbb{N}\cup\{\infty,\omega\}$ the set of all $f\in C^{k}(A,\mathbb{R})$ where $f$ and all of its derivatives up to order $k$ are bounded.
$L^{p}(A,\mathbb{R})$ For $1\leq p<\infty$ the space of $L^{p}$–functions $f:\mathbb{R}^{n}\supseteq A\to\mathbb{R}$, i.e., the space of all (equivalence classes of) measurable functions, for which the integral $\int_{A}|f|^{p}\,\lambda$ converges.
$L^{\infty}(A,\mathbb{R})$ The set of all (equivalence classes of) measurable functions $f$ that are essentially bounded, i.e., for which the essential supremum $\operatornamewithlimits{esup}_{A}|f(x)|<\infty$
@@
$W^{k,\infty}(A,\mathbb{R})$ For $k\in\mathbb{N}$, the space of functions $f\in L^{\infty}(A,\mathbb{R})$ who are weakly kk–times differentiable, and for which all derivatives up to order $k$ are also essentially bounded.
@@
$\mathcal{X}(M)$ The space of all smooth vector fields on $M$.
$\mathcal{X}^{k,\infty}(M)$ For $k\in\mathbb{N}$, the space of vector–fiels who are essentially bounded on $M$, weakly kk–times differentiable, and for which all derivatives up to order $k$ are also essentially bounded. We endow it with the classical $W^{k,\infty}$–norm.
@@ 2.2 Optimal transport
First we recall some well–known results.
@@ 2.2.1 Fundamental results
Let $(M,g)$ be a Riemannian manifold, $x_{0}\in M$, and $1\leq p<\infty$. For a sequence $(\mu_{k})_{k\in\mathbb{N}}$ in $\mathcal{P}_{p}(M)$ and $\mu\in\mathcal{P}_{p}(M)$ the following are equivalent:
@@
    $\mu_{n}\xrightharpoonup{}\mu$ and $\lim_{n\to\infty}\int_{M}d(x,x_{0})^{p}\,\operatorname{d}\mu_{n}(x)=\int_{M}d(x,x_{0})^{p}\,\operatorname{d}\mu(x)$,
  3. 3.
@@
Figalli & Glaudo [7, Theorem 3.1.6] and Lemma 2.2 $\square$
Lemma 2.2.
@@
Figalli & Glaudo [7, Lemma 2.1.13]. $\square$
@@
Let $(M,g)$ be a Riemannian manifold and $\mu,\nu\in\mathcal{W}_{1}(M)$. Then,
$W_{1}(\mu,\nu)=\sup\biggl{\{}\int_{M}f\operatorname{d}\mu-\int_{M}f\operatorname{d}\nu\mid f\in W^{1,\infty}(M),\|f\|_{W^{1,\infty}(M)}\leq 1\biggr{\}}.$ (2.1)
@@
This is a special case of a more general result proved in Villani et al. [23, Theorem 5.10]. $\square$
@@ 2.2.2 Wasserstein embeddings
Let $1\leq q\leq p$. Then the embedding $\iota:\mathcal{W}_{p}(M)\to\mathcal{W}_{q}(M)$ is 11–Lipschitz.
@@
Let $(M,g)$ be a Riemannian manifold, and let $\mu,\nu\in\mathcal{P}_{1}(M)$. Then $t\mapsto t\mu+(1-t)\nu$ is a geodesic in $W^{1}(M)$ connecting $\mu$ and $\nu$.
@@
Let $\mu,\nu\in\mathcal{P}(M)$, and let $\lambda\in[0,1]$. We denote by $\Delta:M\to M\times M$, $x\mapsto(x,x)$. Let $\gamma\in\Gamma^{\operatorname{opt}}(\mu,\nu)$, then $\overline{\gamma}_{\lambda}:=\lambda\Delta_{*}\mu+(1-\lambda)\gamma\in\Gamma(\mu,\lambda\mu+(1-\lambda)\nu)$, since
$(\operatorname{pr}_{1})_{*}(\lambda\Delta_{*}\mu+(1-\lambda)\gamma)$ $=\lambda\mu+(1-\lambda)\mu=\mu$
$(\operatorname{pr}_{2})_{*}(\lambda\Delta_{*}\mu+(1-\lambda)\gamma)$ $=\lambda\mu+(1-\lambda)\nu.$
@@
$W(\mu,\lambda\mu+(1-\lambda)\nu)$ $\leq\int_{M\times M}d(x,y)\,\operatorname{d}\overline{\gamma}_{\lambda}(x,y)$
@@
$=(1-\lambda)W(\mu,\nu)=\lambda W(\mu,\mu)+(1-\lambda)W(\mu,\nu).$
In addition, $W(\lambda\mu+(1-\lambda)\nu,\nu)\leq\lambda W(\mu,\nu)+(1-\lambda)W(\nu,\nu)$, and we have
$W(\mu,\nu)$ $\leq W(\mu,\lambda\mu+(1-\lambda)\nu)+W(\lambda\mu+(1-\lambda)\nu,\nu)$
$\leq(1-\lambda)W(\mu,\nu)+\lambda W(\mu,\nu)=W(\mu,\nu).$
Therefore, $W(\mu,\lambda\mu+(1-\lambda)\nu)=(1-\lambda)W(\mu,\nu)$, so the straight line connects, and the rest follows from the triangle inequality. $\square$

### v2

@@ Abstract
We define new differential structures on the Wasserstein spaces $\mathcal{W}_{p}(M)$ for $p>2$ and a general Riemannian manifold $(M,g)$. We consider a very general and possibly degenerate second order partial differential flow equation with measure dependent coefficients to expand the notion of smooth curves and to ensure that the new differential structure is finer than the classical one. Under weak assumptions, we explicitly construct smooth solutions as uniform limits of Average Flow Approximation Series (a variant of explicit Euler–scheme approximations) in $\mathcal{W}_{p}(M)$ and, thus, prove a generalized version of the Central Limit Theorem. Under slightly stronger assumptions, we prove that smooth solutions of our newly introduced flow–equation are unique.
@@ Acknowledgements
I am grateful to Alessio Figalli and Ulisse Stefanelli for their guidance and encouragement. I would also like to thank Hans Gersbach, Wilhelm Schlag, Darko Mitrovic, Nikolai V. Krylov and Hermann Schichl for their valuable input.
@@ 1 Introduction
Recent years have seen an increased need for analytical calculus on spaces other than finite dimensional vector spaces. In particular, measure spaces like the Wasserstein spaces of optimal transport theory have been of central interest in prominent works like Ambrosio et al. [1, 2], Figalli and Glaudo [9], Santambrogio [27], Villani and others [28], Villani [29].
@@
on the Wasserstein space $\mathcal{W}_{p}(M)$, for a Riemannian manifold $(M,g)$, $p\geq 1$ and some probability vector–field $V$, which remains to be defined. To these means, Otto and Villani [22], Otto [23] were the first to endow the Wasserstein spaces $\mathcal{W}_{p}(\mathbb{R}^{n})$ for $p>1$ with differential structures. The so-called continuity equation
@@
Detailed expositions of the theory in $\mathbb{R}^{n}$ and more general spaces can be found in Ambrosio et al. [1], Gigli [13], Bonnet and Frankowska [6], Cavagnari et al. [7]. However, the so–defined differential structures have important limitations.
@@ 2 Preliminaries
In this preliminary chapter, we introduce some basic notation we shall use in this work and recall some classical results in optimal transport theory and Riemannian geometry. Furthermore, we prove some simple results, which shall be of use later.
2.1 Notation
@@
We will need some real function spaces, especially Sobolev spaces, throughout this work. Let $(Z,d)$ be a complete metric space, $(E,h)$ a Banach manifold and $(M,g)$ a Riemannian manifold. Furthermore, let $A\subset E$, $B\subset Z$ and $N\subset M$ be closed. We denote by $\lambda$ the volume form on $(M,g)$ and define the function spaces (see Michor [19] for formal introductions):
$\mathcal{W}_{p}(Z)$ The Wasserstein space of pp–integrable probability Borel–measures on $M$ for $p\geq 1$ (see [9, 4]).
$\mathcal{W}_{p}(\mu,\nu)$ The pp–Wasserstein distance between $\mu,\nu\in\mathcal{W}_{p}(M)$, for $p\geq 1$ (see [9, 4]).
@@
$\mathcal{P}_{p}(Z)$ The set of all probability Borel–measures on $M$ whose $p$th moment exists for $p\geq 1$.
$\Gamma(\mu,\nu)$ The set of all transport plans between $\mu,\nu\in\mathcal{P}(M)$ (see [9, 4]).
$C(B,\mathbb{R}^{m})$ The set of all continuous functions $f:B\to\mathbb{R}^{m}$.
$C^{k}(A,\mathbb{R}^{m})$ For $k\in\mathbb{N}\cup\{\infty,\omega\}$ the set of all kk–times continuously differentiable functions $f\in C(A,\mathbb{R})$. For $k=\infty$ the set of all smooth functions, and for $k=\omega$ and $A\subset\mathbb{R}^{n}$ the set of all real analytic functions. Also, $C^{0}(A,\mathbb{R}^{m}):=C(A,\mathbb{R}^{m})$.
$C^{k}_{c}(A,\mathbb{R}^{m})$ For $k\in\mathbb{N}\cup\{\infty\}$ the set of all $f\in C^{k}(A,\mathbb{R}^{m})$ with ${\operatorname{supp}}f$ compact.
$C^{k}_{b}(A,\mathbb{R}^{m})$ For $k\in\mathbb{N}\cup\{\infty,\omega\}$ the set of all $f\in C^{k}(A,\mathbb{R}^{m})$ where $f$ and all of its derivatives up to order $k$ are bounded.
$L^{p}(N,\mathbb{R})$ For $1\leq p<\infty$ the space of $L^{p}$–functions $f:M\supseteq N\to\mathbb{R}$, i.e., the space of all (equivalence classes of) measurable functions, for which the integral $\int_{A}|f|^{p}\,\operatorname{d}\lambda$ converges.
$L^{\infty}(A,\mathbb{R})$ The set of all (equivalence classes of) measurable functions $f$ that are essentially bounded, i.e., for which the essential supremum $\operatornamewithlimits{esup}_{A}|f(x)|<\infty$.
@@
$W^{k,\infty}(A,\mathbb{R})$ Let $(E,h)$ be a Banach space or a Riemannian manifold, either with bounded geometry or a fixed atlas. For $k\in\mathbb{N}$, the space of functions $f\in L^{\infty}(A,\mathbb{R})$ who are weakly kk–times differentiable, and for which all derivatives up to order $k$ are also essentially bounded.
@@
$\mathcal{X}(E)$ The space of all smooth vector fields on $E$.
$\mathcal{X}^{k,\infty}(E)$ For $k\in\mathbb{N}$, the space of vector–fiels who are essentially bounded on $E$, weakly kk–times differentiable, and for which all derivatives up to order $k$ are also essentially bounded. We endow it with the classical $W^{k,\infty}$–norm.
@@ 2.2 Optimal transport
First we recall some well–known results of optimal transport in the typical setting of Polish spaces and their less well known extensions to the more general Radon metric spaces. For a comprehensive introduction to optimal transport theory, we refer to Figalli and Glaudo [9] and Villani and others [28].
Recall that a Radon metric space is a metric space $(M,d)$ such that every Borel probability measure on $M$ is a Radon measure.
Definition 2.1 (Radon metric space).
A metric space $(M,d)$ is called a Radon metric space, if every Borel probability measure $\mu$ on $M$ satisfies that for every Borel set $B\subset M$, and every $\varepsilon>0$, there exists a compact set $K_{\varepsilon}\subset B$ such that $\mu(B\setminus K_{\varepsilon})<\varepsilon$.
Most fudamental results of optimal transport theory, which are usually stated for Polish spaces, remain true in Radon metric spaces. This is a strong generalization, since the Radon spaces are far more general objects than Polish spaces. The following result is proved in Fremlin [10, page 244].
Theorem 2.2.
A complete metric space is a Radon metric space if and only if its weigth is measure–free.
The non–existence of metric spaces with measure–free weight is consistent with ZFC—see for example [11]. While it is still an open question whether also the existence of metric spaces with measure–free weight is consistent with ZFC, or whether ZFC implies that all complete metric spaces are Radon metric spaces, all spaces we will consider in this work can be constructed in ZFC and are, therefore, Radon metric spaces.
@@ 2.2.1 Fundamental results
The fundamental result of optimal transport theory is the existence of optimal transport plans between two probability measures.
Theorem 2.3 (Existence of optimal transport plans).
Let $(M,d)$ be a complete Radon metric space and $\mu,\nu\in\mathcal{P}_{p}(M)$. Then, there exists an optimal transport plan $\gamma\in\Gamma(\mu,\nu)$.
Proof.
Bogachev and Kolesnikov [4, Theorem 1.2.1] $\square$
The pp–Wasserstein space is a complete metric space if the underlying metric space is complete.
Lemma 2.4.
Let $(M,d)$ be a complete Radon metric space. Then, $\mathcal{W}_{p}(M)$ is a complete metric space.
Proof.
Mardare et al. [18, Theorem 2.7] $\square$
First, we establish some useful links between the Wasserstein–topology and other classical topologies on spaces of Borel–probability measures. Prokhorov’s theorem characterizes relatively compact sets of probability measures with respect to the narrow topology—see Bogachev [5, Vol. 2 Theorem 8.6.2].
Theorem 2.5 (Prokhorov).
Let $(M,d)$ be a Radon metric space. A set $\mathcal{K}\subseteq\mathcal{P}(M)$ is relatively compact with respect to the narrow topology if and only if it is tight.
Lemma 2.6.
Proof.
Figalli and Glaudo [9, Lemma 2.1.13]. $\square$
Prokhorov’s theorem allows us to directly compare the Wasserstein topology and the narrow toplogy.
@@
Let $(M,d)$ be a Polish metric space, $x_{0}\in M$, and $1\leq p<\infty$. For a sequence $(\mu_{k})_{k\in\mathbb{N}}$ in $\mathcal{P}_{p}(M)$ and $\mu\in\mathcal{P}_{p}(M)$ the following are equivalent:
@@
    $\mu_{n}\xrightharpoonup{}\mu$ and $\lim_{n\to\infty}\int_{M}d(x,x_{0})^{p}\,\operatorname{d}\mu_{n}(x)=\int_{M}d(x,x_{0})^{p}\,\operatorname{d}\mu(x)$.
Furthermore, if $(M,d)$ is locally compact (or in particular a Riemannian manifold), then, the above are also equivalent to
@@
Ambrosio et al. [1, Theorem 7.1.5] $\square$
Corollary 2.8.
Theorem 2.7 also holds if $M$ is any Radon metric space.
@@
Let $(\mu_{n})_{n\in\mathbb{N}}$ be a sequence in $\mathcal{P}_{p}(M)$ and $\mu\in\mathcal{P}_{p}(M)$, such that $\mu_{n}\xrightharpoonup{*}\mu$ and $\lim_{n\to\infty}\int_{M}d(x,x_{0})^{p}\operatorname{d}\mu_{n}(x)=\int_{M}d(x,x_{0})^{p}\operatorname{d}\mu(x)$. Since $M$ is a Radon metric space, there exist compact sets $K_{n,N}\subset M$ and $K_{N}\subset M$, such that for all $n,N\in\mathbb{N}$,
$\mu_{n}(M\setminus K_{n,N})<\frac{1}{N},$ (2.1)
$\mu(M\setminus K_{N})<\frac{1}{N}.$
Define $m\coloneqq\bigcup_{n,N\in\mathbb{N}}K_{n,N}\cup\bigcup_{N\in\mathbb{N}}K_{N}$. In particular, ${\operatorname{supp}}(\mu)\subset\overline{m}$ and for all $n\in\mathbb{N}$, ${\operatorname{supp}}(\mu_{n})\subset\overline{m}$. Since $M$ is a metric space, $\overline{m}$ is separable and complete and, therefore, a Polish space. Now, the result follows from Theorem 2.7 applied to the Polish space $\overline{m}$. $\square$
The next results are some useful tools to determine or bound the Wasserstein distance between two probability measures.
@@
Let $(M,g)$ be a Polish space (or in particular a Riemannian manifold) and $\mu,\nu\in\mathcal{W}_{1}(M)$. Then,
$W_{1}(\mu,\nu)=\sup\biggl\{\int_{M}f\operatorname{d}\mu-\int_{M}f\operatorname{d}\nu\,\biggl|\,f\in{\mathcal{L}ip}(M)\text{, $1$--Lipschitz}\biggr\}.$ (2.2)
@@
This is a special case of a more general result proved in Villani and others [28, Theorem 5.10]. $\square$
If we work on more general metric spaces, we recover a weaker version of the Monge–Kantorovich duality.
Lemma 2.10 (Weak (Monge-)Kantorovich duality).
Let $(M,d)$ be a metric space and $\mu,\nu\in\mathcal{W}_{1}(M)$. Then, for all $f\in{\mathcal{L}ip}(M)$, such that $f$ is 11–Lipschitz,
$\int_{M}f\operatorname{d}\mu-\int_{M}f\operatorname{d}\nu\leq W_{1}(\mu,\nu).$ (2.3)
Proof.
Denote by $\pi_{1},\pi_{2}:M\times M\to M$ the projections on the first and the second coordinates respectively. Let $\gamma$ be a transport plan from $\mu$ to $\nu$. Then,
$\int_{M}f(x)\operatorname{d}\mu(x)-\int_{M}f(y)\operatorname{d}\nu(y)$ (2.4)
$\hskip 18.49988pt=\int_{M}f(\pi_{1}(x,y))\operatorname{d}\gamma(x,y)-\int_{M}f(\pi_{2}(x,y))\operatorname{d}\gamma(x,y)$
$\leq\int_{M}\bigl|f(\pi_{1}(x,y))-f(\pi_{2}(x,y))\bigr|\operatorname{d}\gamma(x,y)$
$\leq\int_{M}d(x,y)\operatorname{d}\gamma(x,y).$
Since $\gamma$ was chosen arbitrarily, the result follows. $\square$
@@ 2.2.2 Wasserstein embeddings
Let $(M,d)$ be a metric space. Let $1\leq q\leq p$. Then the embedding $\iota:\mathcal{W}_{p}(M)\to\mathcal{W}_{q}(M)$ is 11–Lipschitz.
@@
Let $(M,g)$ be a Riemannian manifold, and let $\mu,\nu\in\mathcal{P}_{1}(M)$. Then $t\mapsto t\mu+(1-t)\nu$ is a geodesic in $\mathcal{W}_{1}(M)$ connecting $\mu$ and $\nu$.
@@
Let $\mu,\nu\in\mathcal{P}_{1}(M)$, and let $\lambda\in[0,1]$. We denote by $\Delta:M\to M\times M$, $x\mapsto(x,x)$. Let $\gamma\in\Gamma^{\operatorname{opt}}(\mu,\nu)$, then $\overline{\gamma}_{\lambda}:=\lambda\Delta_{*}\mu+(1-\lambda)\gamma\in\Gamma(\mu,\lambda\mu+(1-\lambda)\nu)$, since
$(\pi_{1})_{*}(\lambda\Delta_{*}\mu+(1-\lambda)\gamma)$ $=\lambda\mu+(1-\lambda)\mu=\mu$
$(\pi_{2})_{*}(\lambda\Delta_{*}\mu+(1-\lambda)\gamma)$ $=\lambda\mu+(1-\lambda)\nu.$
@@
$\mathcal{W}_{1}(\mu,\lambda\mu+(1-\lambda)\nu)$ $\leq\int_{M\times M}d(x,y)\,\operatorname{d}\overline{\gamma}_{\lambda}(x,y)$
@@
$=(1-\lambda)\mathcal{W}_{1}(\mu,\nu)=\lambda\mathcal{W}_{1}(\mu,\mu)+(1-\lambda)\mathcal{W}_{1}(\mu,\nu).$
In addition, $\mathcal{W}_{1}(\lambda\mu+(1-\lambda)\nu,\nu)\leq\lambda\mathcal{W}_{1}(\mu,\nu)+(1-\lambda)\mathcal{W}_{1}(\nu,\nu)$, and we have
$\mathcal{W}_{1}(\mu,\nu)$ $\leq\mathcal{W}_{1}(\mu,\lambda\mu+(1-\lambda)\nu)+\mathcal{W}_{1}(\lambda\mu+(1-\lambda)\nu,\nu)$
$\leq(1-\lambda)\mathcal{W}_{1}(\mu,\nu)+\lambda\mathcal{W}_{1}(\mu,\nu)=\mathcal{W}_{1}(\mu,\nu).$
Therefore, $\mathcal{W}_{1}(\mu,\lambda\mu+(1-\lambda)\nu)=(1-\lambda)\mathcal{W}_{1}(\mu,\nu)$, so the straight line connects, and the rest follows from the triangle inequality. $\square$